# Deepfake-Eval-2024: A Multi-Modal In-the-Wild Benchmark of Deepfakes Circulated in 2024

### v2

@@ Abstract
In the age of increasingly realistic generative AI, robust deepfake detection is essential for mitigating fraud and disinformation. While many deepfake detectors report high accuracy on academic datasets, we show that these academic benchmarks are out of date and not representative of real-world deepfakes. We introduce Deepfake-Eval-2024, a new deepfake detection benchmark consisting of in-the-wild deepfakes collected from social media and deepfake detection platform users in 2024. Deepfake-Eval-2024 consists of 45 hours of videos, 56.5 hours of audio, and 1,975 images, encompassing the latest manipulation technologies. The benchmark contains diverse media content from 88 different websites in 52 different languages. We find that the performance of open-source state-of-the-art deepfake detection models drops precipitously when evaluated on Deepfake-Eval-2024, with AUC decreasing by 50% for video, 48% for audio, and 45% for image models compared to previous benchmarks. We also evaluate commercial deepfake detection models and models finetuned on Deepfake-Eval-2024, and find that they have superior performance to off-the-shelf open-source models, but do not yet reach the accuracy of deepfake forensic analysts. The dataset is available at https://github.com/nuriachandra/Deepfake-Eval-2024.
@@ 1 Introduction
Recent research has shown that people are no longer able to determine whether media is AI-generated or real [7]. Thus, the development of accurate and automated deepfake detection methods has become essential for mitigating the harmful effects of deepfakes. Many deepfake detection models have already been developed, such as GenConViT [8] for video, AASIST [9] for audio, and NPR [10] for image deepfakes, all of which perform extremely well on the academic datasets that they were originally benchmarked on, with AUC values approaching one (Table 5). However, these datasets are not representative of deepfakes circulating on social media, because many use outdated manipulation techniques (e.g., FaceForensics++ [11] and ForgeryNet [12]) with human differentiable fakes (e.g., faces not centered appropriately on heads in ForgeryNet). Further, most existing synthetic datasets have limited content diversity (e.g., exclusively single-scene videos containing a limited number of body poses [13, 14, 15], or exclusively English audio [16, 17, 18]). The most recent in-the-wild deepfake datasets are also outdated, published in 2021 for video [19], and 2022 for audio [18], prior to generative AI advancements such as stable diffusion [20] and the release of commercial models such as ElevenLabs voice conversion [21].
To address the limitations of existing deepfake detection benchmarks, we present Deepfake-Eval-2024, a dataset of deepfakes collected from social media and the free deepfake detection platform TrueMedia.org. Each item in Deepfake-Eval-2024 comes from a social media or TrueMedia.org user who flagged the media as potentially AI-manipulated in 2024. As a result, Deepfake-Eval-2024 is smaller than synthetic datasets, but much more diverse, and directly representative of the deepfakes that individuals encounter in the real world.
We summarize our contributions with the following:
@@ 2 Related Work
Deepfake datasets have not kept up with the fast-moving field of AI content generation. This is particularly true of in-the-wild datasets; the most recent in-the-wild video datasets are from 2020 [22] and 2021 [19]. The only other significant in-the-wild deepfake dataset was audio data published in 2022 [18], which has fewer hours of audio than Deepfake-Eval-2024. Further, to the best of our knowledge, there are no other in-the-wild deepfake image-focused datasets. Supplementary Tables S1, S2, S3 provide a detailed survey of all popular datasets compared to Deepfake-Eval-2024.
The majority of deepfake datasets are synthetically generated, which has enabled the creation of very large datasets (e.g., ForgeryNet with over one million images [12] and ASVspoof datasets [23, 17], Deepfake Detection Challenge (DFDC) [14], DF-Platter [24], and AV-Deepfake1M [25] all with over 300 hours of data). However, due to their synthetic nature, they can fail to capture the characteristics and distribution of deepfakes actually circulated on social media.
@@ 3 Dataset
Table 1: Deepfake-Eval-2024 Video Summary Statistics
Table 2: Deepfake-Eval-2024 Audio Summary Statistics
Real 36.6 1,110 124.80 44.83
Fake 19.9 710 101.51 44.40
All 56.5 1,820 115.46 44.66
Table 3: Deepfake-Eval-2024 Image Summary Statistics
Category Count Mode Resolution
Real 1,208 1,200$\times$1,200
Fake 767 1,024$\times$1,024
All 1,975 1,024$\times$1,024
Deepfake-Eval-2024 is composed of 45 hours of videos and 56.5 hours of audio and 1,975 images. (Tables 1, 2, 3 contain complete summary statistics.) The data includes real, AI-generated, and AI-manipulated content. Audio data includes audio from videos, in addition to audio-only media. The majority of video data has corresponding labeled audio.
An ideal benchmark for deepfake detection is representative of the real-world threat of deepfakes. This requires the following criteria: 1) it contains fake and real content that is difficult for humans to categorize, 2) it includes all popular generative techniques used for deepfake generation, and 3) it has diverse content, representative of the media shared on the internet. Deepfake-Eval-2024 meets all three criteria through its unique data collection approach. All data was collected through users of the deepfake detection platform TrueMedia.org and social media content moderation forums.
@@ 3.1 Data Collection
Data Sources. The deepfake detection platform TrueMedia.org was a non-profit application originally used primarily by journalists and fact-checkers starting in April 2024 before becoming available to the general public in September 2024. Deepfake-Eval-2024 includes data uploaded to TrueMedia.org. Users provided a social media link or directly uploaded content to be checked for AI-manipulation. We also created a bot on X (previously known as Twitter) that allowed users to add content to our platform by tagging the bot. In addition, we uploaded posts from X that had been flagged by X Community Notes as potentially manipulated media. The top five most common data sources of Deepfake-Eval-2024 are X, direct upload, TikTok, Instagram, and Youtube (Figure 2, Figure S1). Direct upload indicates that the media was uploaded directly to the TrueMedia.org deepfake detection platform rather than a social media or website link.
Figure 2: Origins of data in Deepfake-Eval-2024 separated by modality. In total, media was shared from 88 different web-domain names.
Collection Ethics. TrueMedia.org users were informed in the Terms of Use that ‘anything you share will not be private.’ We are not aware of any sensitive personal identification information included in the dataset. However, as with any media sourced from the internet, the data could contain information about an individual that they did not expressly consent to share.
@@
Dataset Diversity. Collecting data through social media and deepfake detection platform users also provides increased diversity with respect to generative models, ethnicity, language, and content. Deepfake-Eval-2024 is a sample of currently circulating AI-generated content. Thus, we estimate that our dataset includes AI-generated and manipulated content from every type of contemporary model commonly used to generate deepfakes. Further, TrueMedia.org users came from all over the world, resulting in increased ethnic and linguistic diversity in our dataset. Our dataset is 78.7% English and includes a total of 52 different languages (Figure 3). The content of the media itself has larger variations than the clean standardized content typically found in academic datasets. The diversity in data origins (Figure 2) results in our dataset containing a wide variety of different media styles, including videos of political speeches and self-shot content-creators, images of large crowds and close-up portraits, and audio clips of debating politicians and background music.
Figure 3: Language distribution in Deepfake-Eval-2024 audio and video data. The dataset contains a total of 52 different languages (42 languages in Deepfake-Eval-2024-audio and 49 languages in Deepfake-Eval-2024-video). Languages were identified using speech recognition model Whisper, which is known to perform well on language identification tasks [30].
@@ 3.2 Data Filtering
We remove duplicate data using a combination of manual review and hash functions. We include cases where two pieces of media have minor variations and thus appear to be the same (e.g., different cropping of the same video).
In order to tailor our datasets to evaluate deepfake detection models, we remove images and videos that do not contain photorealistic faces. This resulted in the removal of cartoons, art, and scenes without humans. We use GPT-4o (version 2024-08-06) to identify images with photorealistic faces. We note that GPT-4o’s responses have high precision but also a high false negative rate. To account for this, we manually review all images marked as non-photorealistic faces by GPT-4o. To identify videos with photorealistic faces we first use the dlib face detection library [31] to determine whether each frame contains a face or not. Videos where no faces are detected are then reviewed manually to check for missed faces.
@@ 3.3 Data Labeling
Label Definition. We label media as fake if it was AI-generated or manipulated. We choose to define our labels this way despite the challenge of differentiating AI-manipulated content from traditional forms of manipulation so that this dataset can be used to benchmark deepfake detection models, the majority of which are trained to differentiate between AI-generated and real content.
@@
Deepfake Forensic Analysis. We rely on articles published by professional fact-checking organizations such as Snopes [^1] Snopes is one of the oldest and most widely trusted fact-checking websites, operating since 1994. The site employs professional journalists who investigate viral claims and urban legends. https://www.snopes.com/about/ and AFP Fact Check [^2] AFP Fact Check is the fact-checking division of Agence France-Presse, a major world news agency. They operate in multiple languages and verify content across various social media platforms. https://factcheck.afp.com/about-afp to confirm if AI manipulation was used. Additionally, we utilize community moderation platforms like X Community Notes to locate and review primary sources. We conduct source context verification by comparing social media posts with original materials such as full-length videos to verify if media has been manipulated. We further scrutinize the media for evidence of AI manipulation using specific forensic markers. For example, we rely on the synchronization of mouth movements with vocal sounds as a primary measure of authenticity in video and audio media. For videos and images, face-swaps are classified as fake if they were created after 2023, and unknown otherwise. This time separation is supported by research that reported a 704% increase in AI-generated face swaps during 2023 [32], suggesting that the majority of face-swaps created in 2023 and later are done so using AI. Other common indicators of AI manipulation, including anatomical implausibilities, sociocultural implausibilities, and stylistic artifacts, are identified based on the framework created by Kamali et al. [33]. When the label of audio media cannot be determined from sources, due to the challenge of manually detecting AI-generated audio (and differentiating it from non-AI generated voice impersonators), media is marked as fake if and only if there are both audible traits indicating that it is fake (e.g. sociocultural implausibilities), in addition to at least two commercial audio detectors predicting the media as fake. We note that this labeling approach results in audio labels that are more likely to be correlated with existing detectors, which may result in a less challenging audio dataset. For all modalities, when labelers are uncertain about the presence or absence of forensic markers, they label the media as “unknown.” See Appendix B for complete taxonomy and media examples.
@@
Labeling Accuracy. To assess the consistency and quality of our annotations, the labeling team lead checked a random sample of 10% of the data for each of the three modalities. Annotations created by the team lead were excluded to avoid self-assessment bias. For videos, we find a 6.6% disagreement between labelers, with the largest discrepancy between real and unknown at 3.7%. For audio, we find a 7.9% disagreement between labelers, with the largest discrepancies between real and unknown at 3.5% and between fake and unknown at 3.8%. And for images, we find a 9% disagreement between labelers, with the largest discrepancy between fake and unknown at 6%. (See Table 4 for complete disagreement breakdown.) Given that the inter-labeler disagreement was consistently below 10%, we posit that deepfake detection models should be able to achieve at least 90% accuracy on Deepfake-Eval-2024, and likely higher given that real vs. fake disagreement is below 2.5%.
Table 4: Inter-rater Disagreement Statistics Across Modalities
@@
Common human errors include: differentiating between dubbed videos (where the video has not been AI-manipulated, and thus is real), and lipsynced videos (where the video has been AI-manipulated to make the mouth match new words and thus should be marked as fake); determining if audio sound is synthetic; and missing anatomical or sociocultural implausibilities.
@@ 4 Experiments
Commercial Models. We evaluate commercially available deepfake detection models from companies that partnered with TrueMedia.org: Hive, Reality Defender, Pindrop, AI or Not, Hiya, Fraunhofer, and Sensity AI. Many companies provide multiple models. In total, we evaluate 22 different commercial models (six video, eight audio, and eight image models). We anonymize the names of the models and providers to comply with contractual agreements. The commercial models are evaluated using their latest available versions as of December 2024. All vendors were blind to the test data. Due to the high per-query cost of commercial vendors, we were unable to evaluate all commercial models on the entirety of Deepfake-Eval-2024. Instead, we evaluate all models on a subset of Deepfake-Eval-2024, and then evaluate the top two models for each modality on the entire Deepfake-Eval-2024. We report the performance of the top commercial model for each modality in Table 7.
Evaluation Metrics. The field of deepfake detection does not have a standard set of evaluation metrics, and past publications have selectively focused on specific metrics (e.g. AUC in [41], accuracy in [35], and EER in [29]). Selective reporting on specific metrics makes comparison across publications challenging and does not offer a comprehensive view of model performance. As such, we evaluate the performance of models on Deepfake-Eval-2024 through a wide variety of metrics focusing on AUC, precision, recall, and F1-score, with full metrics including accuracy, false positive rate, and false negative rate available in the supplementary materials. We also report EER for audio models in the supplementary, as it is common in audio deepfake literature. Some open-source and commercial models fail to run on all media files due to model constraints (e.g., media length limits, or requirements for a face to be detected in a certain number of frames). When a model fails to produce a prediction, we exclude this file when calculating the metrics for the associated model.
@@
Finetuning. In order to determine if models can improve real-world deepfake detection performance by training on more representative data, we finetune all open-source models on 60% of Deepfake-Eval-2024, and evaluate the performance on the remaining 40% of the data (Table 6). This split mirrors real-world scenarios where models must generalize from limited training data to detect unseen deepfake techniques. We finetune each model following the original authors’ recommended training procedures and hyperparameters where available, using early stopping to avoid overfitting.
Table 5: Open-Source Model Performance on Deepfake-Eval-2024
@@
Original publication test data includes the following datasets for each model. Where multiple datasets are specified, the reported metrics are averages over these datasets. GenConViT: [11], [14], [42], [43]; FTCN: [42]; Styleflow: [42], [44], [45], [46]; AASIST, RawNet2, and P3 were all evaluated on the LA eval set of ASVspoof2019 [23]; UFD: [47] and subsets of LAION-400M [48] and AI generated images from latent diffusion models [20], Glide [49], and DALL·E mini [50] provided by the original publication [35]; DistilDIRE: ImageNet and AI generated images from Stable Diffusion v1 [20] and ADM [51] as specified in the original publication [36]; NPR: [52], [27], and the dataset from [35].

### v4

@@ Abstract
In the age of increasingly realistic generative AI, robust deepfake detection is essential for mitigating fraud and disinformation. While many deepfake detectors report high accuracy on academic datasets, we show that these academic benchmarks are out of date and not representative of real-world deepfakes. We introduce Deepfake-Eval-2024, a new deepfake detection benchmark consisting of in-the-wild deepfakes collected from social media and deepfake detection platform users in 2024. Deepfake-Eval-2024 consists of 45 hours of videos, 56.5 hours of audio, and 1,975 images, encompassing the latest manipulation technologies. The benchmark contains diverse media content from 88 different websites in 52 different languages. We find that the performance of open-source state-of-the-art deepfake detection models drops precipitously when evaluated on Deepfake-Eval-2024, with AUC decreasing by 50% for video, 48% for audio, and 45% for image models compared to previous benchmarks. We also evaluate commercial deepfake detection models and models finetuned on Deepfake-Eval-2024, and find that they have superior performance to off-the-shelf open-source models, but do not yet reach the accuracy of deepfake forensic analysts.
@@ 1 Introduction
Recent research has shown that people are no longer able to determine whether media is AI-generated or real [7]. Thus, the development of accurate and automated deepfake detection methods has become essential for mitigating the harmful effects of deepfakes. Many deepfake detection models have already been developed, such as GenConViT [8] for video, AASIST [9] for audio, and NPR [10] for image deepfakes, all of which perform extremely well on the academic datasets that they were originally benchmarked on, with AUC values approaching one (Table 3). However, these datasets are not representative of deepfakes circulating on social media, because many use outdated manipulation techniques (e.g., FaceForensics++ [11] and ForgeryNet [12]) with human differentiable fakes (e.g., faces not centered on heads in ForgeryNet). Further, most existing synthetic datasets have limited content diversity (e.g., exclusively single-scene videos containing a limited number of body poses [13, 14, 15], or exclusively English audio [16, 17, 18]). The most recent in-the-wild deepfake datasets are also outdated, published in 2021 for video [19], and 2022 for audio [18], prior to generative AI advancements such as stable diffusion [20] and the release of commercial models such as ElevenLabs voice conversion [21].
To address the limitations of existing deepfake detection benchmarks, we present Deepfake-Eval-2024, a dataset collected from social media and the free deepfake detection platform TrueMedia.org. Each item in Deepfake-Eval-2024 comes from a social media or TrueMedia.org user who flagged the media as potentially AI-manipulated in 2024. As a result, Deepfake-Eval-2024 is smaller than synthetic datasets, but much more diverse, and directly representative of the deepfakes that individuals encounter in the real world. We summarize our contributions with the following:
@@ 2 Related Work
Deepfake datasets have not kept up with the fast-moving field of AI content generation. This is particularly true of in-the-wild datasets; the most recent in-the-wild video datasets are from 2020 [22] and 2021 [19]. The only other in-the-wild audio deepfake dataset was published in 2022 [18], and has fewer hours of audio than Deepfake-Eval-2024. Further, to the best of our knowledge, there are no other in-the-wild image-focused deepfake datasets. Supplementary Tables S1, S2, S3 provide a detailed survey of popular datasets compared to Deepfake-Eval-2024.
The majority of deepfake datasets are synthetically generated, which has enabled the creation of very large datasets (e.g., ForgeryNet with over one million images [12] and ASVspoof datasets [23, 17], Deepfake Detection Challenge (DFDC) [14], DF-Platter [24], and AV-Deepfake1M [25] all with over 300 hours of data). However, due to their synthetic nature, they can fail to capture the characteristics and distribution of deepfakes circulated on social media.
@@ 3 Dataset
Table 1: Deepfake-Eval-2024 Summary Statistics
Modality Size Avg Resolution Video 45.1 hrs 2,036 FPS, 576$\times$720 px Audio 56.5 hrs 44.66 kHz Image 1,975 images 1,024$\times$1,024 px
Deepfake-Eval-2024 is composed of 45 hours of videos and 56.5 hours of audio and 1,975 images. (Table 1 contains abbreviated summary statistics, with complete statistics stratified by label available in Supplementary Tables S4, S5, and S6.) The data includes real, AI-generated, and AI-manipulated content. Audio data includes audio from videos, in addition to audio-only media. The majority of video data has corresponding labeled audio.
An ideal benchmark for deepfake detection is representative of the real-world threat of deepfakes. This requires the following criteria: 1) it contains fake and real content that is difficult for humans to categorize, 2) it includes all popular generative techniques used for deepfake generation, and 3) it has diverse content, representative of media shared on the internet. We will show that Deepfake-Eval-2024 meets all three criteria through its unique data collection approach. All data was collected through the deepfake detection platform TrueMedia.org and social media content moderation forums.
@@ 3.1 Data Collection
Data Sources. The deepfake detection platform TrueMedia.org was a non-profit application used by journalists and fact-checkers starting in April 2024, and used by the general public starting in September 2024. Deepfake-Eval-2024 includes data uploaded to TrueMedia.org. Users provided a social media link or directly uploaded content to be checked for AI-manipulation. We also created a bot on X (previously known as Twitter) that allowed users to add content to our platform by tagging the bot. The code for this bot is available at https://github.com/truemediaorg/socialbot. In addition, we uploaded X posts that had been flagged by X Community Notes as potentially manipulated media. The top five most common data sources of Deepfake-Eval-2024 are X, direct upload, TikTok, Instagram, and Youtube (Figure 2, Supp. Figure S1).
Figure 2: Distribution of data origins and languages in Deepfake-Eval-2024. In total, media was shared from 88 different web-domain names. The dataset also contains a total of 52 different languages (42 languages in Deepfake-Eval-2024-audio and 49 languages in Deepfake-Eval-2024-video). Languages were identified using speech recognition model Whisper [30].
Collection Ethics. TrueMedia.org users were informed in the Terms of Use that ‘by sharing your content, you agree to follow our content terms and you agree everything you share is consistent with these terms. Anything you share will not be private and can be used by us, our partners, and others we work with, in lots of different ways.’ To the best of our knowledge the dataset does not contain any personal identification information (i.e. social security numbers, emails, financial information, or medical records), and due to the nature of our data (non-textual), it is highly unlikely that we have incidentally collected such information through either TrueMedia.org or other web-sources.
@@
Dataset Diversity. Collecting data through social media and deepfake detection platform users also provides increased diversity with respect to generative models, ethnicity, language, and content. Deepfake-Eval-2024 is a sample of currently circulating AI-generated content. Thus, we estimate that our dataset includes AI-generated and manipulated content from every type of commonly used contemporary model deepfake generation model. Further, TrueMedia.org users came from all over the world, resulting in increased ethnic and linguistic diversity in our dataset. Our dataset is 78.7% English and includes a total of 52 different languages (Figure 2). The content of the media itself has larger variations than the standardized content typically found in academic datasets. The diversity in data origins (Figure 2, Supp. Figure S1) results in a wide variety of different media styles, including videos of political speeches and self-shot content-creators, images of large crowds and close-up portraits, and audio clips of debating politicians.
@@ 3.2 Data Filtering
We remove duplicate data using a combination of manual review and hash functions. We include cases where two pieces of media have minor variations and thus appear to be the same (e.g., different cropping of the same video). In order to tailor our datasets to evaluate deepfake detection models, we remove images and videos that do not contain photorealistic faces. This resulted in the removal of cartoons, art, and scenes without humans. We use GPT-4o (version 2024-08-06) to identify images with photorealistic faces. We note that GPT-4o’s responses have high precision but also a high false negative rate. To account for this, we manually review all images marked as non-photorealistic faces by GPT-4o. To identify videos with photorealistic faces we first use the dlib face detection library [31] to determine whether each frame contains a face or not. Videos where no faces are detected are then reviewed manually to check for missed faces.
@@ 3.3 Data Labeling
Table 2: Inter-rater disagreement statistics
Label Definition. We label media as fake if it was AI-generated or manipulated. We choose to define our labels this way despite the challenge of differentiating AI-manipulated content from traditional forms of manipulation so that this dataset can be used to benchmark deepfake detection models which are trained to differentiate between AI-generated and real content.
@@
Deepfake Forensic Analysis. We rely on articles published by professional fact-checking organizations such as Snopes [32] and AFP Fact Check [33] to confirm if AI manipulation was used. Additionally, we utilize community moderation platforms like X Community Notes to locate and review primary sources. We conduct source context verification by comparing social media posts with original materials such as full-length videos to verify if media has been manipulated. We further scrutinize the media for evidence of AI manipulation using specific forensic markers. For example, we rely on the synchronization of mouth movements with vocal sounds as a primary measure of authenticity in video and audio media. For videos and images, face-swaps are classified as fake if they were created after 2023, and unknown otherwise. This time separation was chosen based on research that suggests that the majority of face-swaps created in 2023 or later use AI [34]. Other common indicators of AI manipulation, including anatomical implausibilities, sociocultural implausibilities, and stylistic artifacts, are identified based on the framework created by Kamali et al. [35]. When the label of audio media cannot be determined from sources, due to the challenge of differentiating AI generated audio from non-AI generated voice impersonators, media is marked as fake if and only if there are both audible traits indicating that it is fake (e.g. sociocultural implausibilities), in addition to at least two commercial audio detectors predicting the media as fake. Audio detectors alone are never used to label. We note that this labeling approach results in audio labels that are correlated with existing detectors, which is a limitation of the audio dataset. However, this approach is necessary, as it is impossible to differentiate between socioculturally-implausible audio made by highly skilled impersonators (real audio) and AI-generated audio (see Jordan Peele’s highly accurate Barack Obama impersonation which uses real audio and AI-manipulated video as an example [36, 37]). See Appendix B for complete labeling taxonomy and media examples.
@@
Labeling Accuracy. To assess the consistency and quality of our annotations, the labeling team lead double-reviewed a random sample of 10% of the data. Annotations created by the team lead were excluded to avoid self-assessment bias. For videos, we find a 6.6% disagreement between labelers, with the largest discrepancy between real and unknown at 3.7%. For audio, we find a 7.9% disagreement between labelers, with the largest discrepancies between real and unknown at 3.5% and between fake and unknown at 3.8%. And for images, we find a 9% disagreement between labelers, with the largest discrepancy between fake and unknown at 6%. (See Table 2 for complete disagreement breakdown.) This disagreement between trained lablers represents the challenge of the task. Given that the inter-labeler disagreement was consistently below 10% , we posit that deepfake detection models should be able to achieve at least 90% accuracy on Deepfake-Eval-2024, and likely higher given that real vs. fake disagreement is always below 2.5%.
@@
Common labeling errors include: differentiating between dubbed videos (where the video has not been AI-manipulated, and thus is real), and lipsynced videos (where the video has been AI-manipulated to make the mouth match new words and thus should be marked as fake); determining if audio sound is synthetic; and missing anatomical or sociocultural implausibilities.
@@ 4 Experiments
Commercial Models. We evaluate commercially available deepfake detection models from companies that partnered with TrueMedia.org: Hive, Reality Defender, Pindrop, AI or Not, Hiya, Fraunhofer, and Sensity AI. Many companies provide multiple models. In total, we evaluate 22 different commercial models (six video, eight audio, and eight image models). We anonymize the performance of the models and providers to comply with contractual agreements. Commercial models are evaluated using their latest available versions as of December 2024. All vendors were blind to the test data. Due to the high per-query cost of commercial vendors, we were unable to evaluate all commercial models on the entirety of Deepfake-Eval-2024. Instead, we evaluate all models on a subset of Deepfake-Eval-2024, and then evaluate the top two models for each modality on the entire Deepfake-Eval-2024. We report the performance of the top commercial model for each modality in Table 5.
Evaluation Metrics. Past deepfake detection publications have selectively focused on specific metrics (e.g. AUC in [45], accuracy in [39], and EER in [29]). Selective reporting on specific metrics makes comparison across publications challenging and does not offer a comprehensive view of model performance. As such, we evaluate the performance of models on Deepfake-Eval-2024 through a wide variety of metrics focusing on AUC, F1-score, and accuracy, with full metrics including precision, recall, false positive rate, and false negative rate available in the supplementary materials. We also report EER for audio models in the supplementary, as it is common in audio deepfake literature. Some open-source and commercial models fail to run on all media files due to model constraints (e.g., media length limits, or requirements for a face to be detected in a certain number of frames). When a model fails to produce a prediction, we exclude this file when calculating the metrics for the associated model.
@@
Finetuning. To determine if real-world deepfake detection performance can be improved by training on representative data, we finetune all open-source models on 60% of Deepfake-Eval-2024, and evaluate the performance on the remaining 40% of the data (Table 4). This split mirrors real-world scenarios where models must generalize from limited training data to detect unseen deepfake techniques. We finetune each model following the original authors’ recommended training procedures and hyperparameters where available, using early stopping to avoid overfitting.
Table 3: Open-source model performance on Deepfake-Eval-2024 and original benchmarks
@@
Original publication test data includes the following datasets for each model. Where multiple datasets are specified, the reported metrics are averages over these datasets. GenConViT: [11], [14], [46], [47]; FTCN: [46]; Styleflow: [46], [48], [49], [50]; AASIST, RawNet2, and P3 were all evaluated on the LA eval set of ASVspoof2019 [23]; UFD: [51] and subsets of LAION-400M [52] and AI generated images from latent diffusion models [20], Glide [53], and DALL·E mini [54] provided by the original publication [39]; DistilDIRE: ImageNet and AI generated images from Stable Diffusion v1 [20] and ADM [55] as specified in the original publication [40]; NPR: [56], [27], and the dataset from [39]. Complete performance metrics on Deepfake-Eval-2024 are available at Supp. Table S8.